# SCILLA: SurfaCe Implicit Learning for Large Urban Area, a volumetric hybrid solution

### v2

@@ Appendix A Additional implementation details
We report in table 4 the hash grid encoding parameters from Instant-NGP [23]. We summarize the split of KITTI-360 [18] sequences used for our evaluations in table 5. We use all four cameras for KITTI-360 [18], and the three front cameras for Pandaset [38], nuScenes [3] and Waymo Open Dataset  [29].
@@
Seq. Start End # frames per cam.
30 1728 1822 48
31 2890 2996 54
35 980 1092 57
36 112 166 28
Appendix B Additional results
nuScenes [3] Waymo Open Dataset [29]
RGB Refer to caption Refer to caption Refer to caption Refer to caption
StreetSurf [11] Refer to caption Refer to caption Refer to caption Refer to caption
(a) Seq. 34 (b) Seq. 664 (c) Seq. 13196 (d) Seq. 14869
Figure 9: Qualitative experiments on both nuScenes [3] (Left), and Waymo [29] (Right). We report the RGB ground-truth image. In a different point-of-view than the GT we compare our generated SDF mesh to StreetSurf’s mesh.
B.1 Additional qualitative results
We report in figure 9 additional qualitative results on nuScenes [3] and Waymo Open Dataset [29]. We find that SCILLA reconstructs higher-quality surfaces compared to StreetSurf and can recover many scene details (see red-boxes highlighted on the figures).
B.2 Ablation study
We ablate the effect of random sample attribution and our proposed probability-based samples attribution. The qualitative results at various training steps are presented in figure 10. The results demonstrate that SCILLA samples attribution strategy initially learns the coarse geometry of the scene during early training stages. While random sample attribution can approximate an accurate SDF representation by the end of the hybrid stage, it results in an incomplete mesh compared to the mesh generated using probability-based sample attribution.
@@ B.2 Ablation study
Appendix C Applications
C.1 Tanks And Temples dataset
In this section, we report SCILLA’s qualitative results on Tanks and Temples dataset [16]. The results reported in figure 11 demonstrate that SCILLA is not restricted to driving datasets and urban areas. SCILLA can represent large, detailed, indoor and outdoor sequences from Tanks And Temples. We report the results of SCILLA with and without the hybrid volumetric stages. Our observation suggests that the transition from the volumetric to surface representation enables to recover various scene details. In addition, SCILLA’s training time takes less than an hour, compared to other state-of-the-art indoor and outdoor landmarks reconstruction methods, such as Neuralengelo [17] which requires 10h training time.
@@ C.1 Tanks And Temples dataset
C.2 Textured mesh
Due to the high-quality of SCILLA’s reconstructed surfaces, we can leverage modern Multi-View Stereo (MVS) tools like OpenMVS [4] to produce detailed and colorized representations of driving sequences. We find that SCILLA’s textured mesh is more complete and accurate compared to StreetSurf’s textured mesh.
@@ C.2 Textured mesh
Figure 12: We use OpenMVS [4] to colorize the outputted meshes. We compare our colored mesh to StreetSurf [11], for the sequence 102751 from Waymo Open Dataset [29].
Refer to caption Refer to caption Refer to caption Refer to caption
(a) RGB (b) Predicted sky mask (c) SCILLA (ours) (d) StreetSurf [11]
@@
Refer to caption Refer to caption Refer to caption Refer to caption
(a) RGB (b) RGB highlighted (c) SCILLA (ours) (d) StreetSurf [11]
@@
Refer to caption Refer to caption Refer to caption Refer to caption
(a) RGB (b) Monocular normal (c) SCILLA (ours) (d) StreetSurf [11]
@@ Appendix D Limitations
  • Disentangling fine details from the sky: unlike StreetSurf [11], which models close range, far range, and sky separately, SCILLA separates only the sky from the other scene’s modeling. However, SCILLA may struggle to distinguish fine details from the sky, particularly in cases where objects are thin, as shown in figure 13.
  • Fine details in wide sequences: SCILLA sometimes fails in accurately reconstructing scene details in wide and open sequences, such as scene 23 from Pandaset reported in figure 14.
  • Inaccurate road reconstruction: for sequence 664 from nuScenes (see figure 15), SCILLA faces challenges in reconstructing the road due to inaccuracies in the monocular normal prediction from Omnidata [8]. In comparison, StreetSurf [11] demonstrates more accurate road reconstruction, attributed to its road-surface initialization.

### v3

@@ 9 Additional implementation details
We report in table 4 the hash grid encoding parameters from Instant-NGP [25]. We summarize the split of KITTI-360 [20] sequences used for our evaluations in table 5. We use all four cameras for KITTI-360 [20], and the three front cameras for Pandaset [41], nuScenes [3] and Waymo Open Dataset [32]. We sample one image out of two for KITTI-360, and one image out of 8 for the other datasets.
@@
Seq. KITTI Sync. Start End # frames per cam.
30 0004 1728 1822 48
31 0009 2890 2996 54
35 0009 980 1092 57
36 0010 112 166 28
10 Additional results
10.1 Additional qualitative results
RGB (ref. only) COLMAP [30] OpenMVS [4] StreetSurf [11] SCILLA (ours)
Seq. 35 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(a) KITTI-360 [20]
RGB (ref. only) COLMAP [30] OpenMVS [4] StreetSurf [11] SCILLA (ours)
Seq. 37 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(b) Pandaset [41]
RGB (ref. only) COLMAP [30] OpenMVS [4] StreetSurf [11] SCILLA (ours)
Seq. 34 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Seq. 664 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(c) nuScenes [3]
RGB (ref. only) COLMAP [30] OpenMVS [4] StreetSurf [11] SCILLA (ours)
Seq. 1319 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Seq. 1486 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(d) Waymo Open Dataset [32]
Figure 8: Qualitative experiments results on 8(a) KITTI-360s [20], 8(b) Pandaset [41], 8(c) nuScenes [3] and 8(d) Waymo Open Dataset [32]. RGB ground-truth images shown for reference of the scene layout. Complex scene geometries reconstructed in the mesh are highlighted in red squares. We compare our generated SDF mesh to COLMAP, OpenMVS and StreetSurf [11] meshes.
We report in Fig.8 additional qualitative results on nuScenes [3] and Waymo Open Dataset [32]. We find that SCILLA reconstructs higher-quality surfaces compared to StreetSurf and can recover many scene details (see highlighted red-boxes on the figures).
10.2 Ablation study
We ablate the effect of random sample attribution and our proposed probability-based samples attribution. The qualitative results at various training steps are presented in  Fig.9. The results demonstrate that SCILLA samples attribution strategy initially learns the coarse geometry of the scene during early training stages. While random sample attribution can approximate an accurate SDF representation by the end of the hybrid stage, it results in an incomplete mesh compared to the mesh generated using probability-based sample attribution.
@@ 10.2 Ablation study
11 Applications
11.1 Tanks And Temples dataset
In this section, we report SCILLA’s qualitative results on Tanks and Temples dataset [18]. The results reported in Fig.  10 demonstrate that SCILLA is not restricted to driving datasets and urban areas. SCILLA can represent large, detailed, indoor and outdoor sequences from Tanks And Temples. We report the results of SCILLA with and without the hybrid volumetric stages. Our observation suggests that the transition from the volumetric to surface representation enables to recover various scene details. In addition, SCILLA’s training time takes less than an hour, compared to other state-of-the-art indoor and outdoor landmarks reconstruction methods, such as Neuralengelo [19] which requires 24h of training.
@@ 11.1 Tanks And Temples dataset
11.2 Textured mesh
Due to the high-quality of SCILLA’s reconstructed surfaces, we can leverage modern Multi-View Stereo (MVS) tools like OpenMVS [4] to produce detailed and colorized representations of driving sequences. As shown in Fig. 11, we find that SCILLA’s textured mesh is more complete and accurate compared to StreetSurf’s textured mesh.
@@ 11.2 Textured mesh
Figure 11: We use OpenMVS [4] to assign texture to the outputted meshes. We compare our colored mesh to StreetSurf [11], for the sequence 102751 from Waymo Open Dataset [32].
Refer to caption Refer to caption
(a) RGB (b) Predicted sky mask
Refer to caption Refer to caption
(c) SCILLA (ours) (d) StreetSurf [11]
@@
Refer to caption Refer to caption
(a) RGB (b) RGB highlighted
Refer to caption Refer to caption
(c) SCILLA (ours) (d) StreetSurf [11]
@@
Refer to caption Refer to caption
(a) RGB (b) Monocular normal
Refer to caption Refer to caption
(c) SCILLA (ours) (d) StreetSurf [11]
@@ 12 Limitations
  • Disentangling fine details from the sky: unlike StreetSurf [11], which models close range, far range, and sky separately, SCILLA separates only the sky from the other scene’s modeling. However, SCILLA may struggle to distinguish fine details from the sky, particularly in cases where objects are thin, as shown in Fig. 12.
  • Fine details in wide sequences: SCILLA sometimes fails in accurately reconstructing scene details in wide and open sequences, such as scene 23 from Pandaset reported in Fig. 13.
  • Inaccurate road reconstruction: for sequence 664 from nuScenes (see Fig. 14), SCILLA faces challenges in reconstructing the road due to inaccuracies in the monocular normal prediction from Omnidata [8]. In comparison, StreetSurf [11] demonstrates more accurate road reconstruction, attributed to its road-surface initialization.